# StorySparkQA: Expert-Annotated QA Pairs with Real-World Knowledge for Children’s Story-Based Learning

### v2

@@ Abstract
Interactive story reading is a common parent-child activity, where parents expect to teach both language skills and real-world knowledge beyond the story. While increasing storytelling and reading systems have been developed for this activity, they often fail to infuse real-world knowledge into the conversation. This limitation can be attributed to the existing question-answering (QA) datasets used for children’s education, upon which the systems are built, failing to capture the nuances of how education experts think when conducting interactive story reading activities. To bridge this gap, we design an annotation framework, empowered by existing knowledge graph to capture experts’ annotations and thinking process, and leverage this framework to construct StorySparkQA dataset, which comprises $5,868$ expert-annotated QA pairs with real-world knowledge. We conduct automated and human expert evaluations across various QA pair generation settings to demonstrate that our StorySparkQA can effectively support models in generating QA pairs that target real-world knowledge beyond story content. StorySparkQA  [^1] The StorySparkQA dataset and annotation framework can also be found in: https://github.com/neuhai/StorySparkQA is available at https://huggingface.co/datasets/NEU-HAI/StorySparkQA.
@@
Jiaju Chen East China Normal University                        Yuxuan Lu Northeastern University                        Shao Zhangthanks: Work done when visiting Northeastern University. Shanghai Jiao Tong University
@@
Yuling Sunthanks:  Corresponding Author: ylsun@cs.ecnu.edu.cn. East China Normal University
@@ 1 Introduction
Interactive story reading is a common parent-child activity, where parents often sit together with preschool children, read storybooks, and proactively engage in question-answering (QA) conversations with them Wright (1995); Isbell et al. (2004). Such guided conversations are typically grounded in but beyond the story narratives Kotaman (2013), with parents’ expectations of guiding children to learn real-world knowledge and improving their historical, cultural, and emotional awareness Sun et al. (2024). This immersive story-based interaction has been proven to be effective in better supporting preschooler knowledge learning Zhang et al. (2024), enhancing their reading comprehension capabilities Xu et al. (2021), etc.
Despite the benefits of interactive story reading, parents often struggle to appropriately conduct such interactive story reading with children because of multi-facet difficulties Golinkoff et al. (2019); Sun et al. (2024). Specifically, such interactive story reading needs parents to identify the knowledge of interest during storytelling, formulate the real-world knowledge piece they want to teach in mind ("what to ask"), then ask an engaging question ("how to ask") to children at the appropriate time ("when to ask"). Yet, most parents lack the educational expertise necessary to guide such educational conversations Golinkoff et al. (2019); Sun et al. (2024). Also, parents in contemporary society often hardly maintain high concentration to accompany their children due to the need to deal with other work and family chores at the same time Zhang et al. (2022); Sun et al. (2024).
Recently, AI-assisted storytelling systems (e.g. StoryBuddy Zhang et al. (2022), TaleMate Vargas-Diaz et al. (2023), MatheMyths Zhang et al. (2024)), have demonstrated effectiveness in children’s storytelling scenarios Dietz et al. (2021). These systems primarily take advantage of the verbal communication interface and advanced language models to support natural conversation with humans Mahmood et al. (2023); Chan et al. (2024); Yang et al. (2024); Dietz et al. (2021). Nevertheless, existing AI-assisted storytelling and reading systems are not without limitations. Particularly, building on top of data resources with mostly extractive QA pairs (e.g., FairytaleQA Xu et al. (2022)) – where the answers can be found directly in the story narrative – these systems fall short in helping parents teach real-world knowledge beyond the story narrative Yao et al. (2021), which is one of the main expectations of parents Sun et al. (2024).
@@
StorySparkQA can benefit different research directions in the children education domain, particularly in better understanding domain experts’ thinking process, and training models to generate story-based QA pairs infused with real-world knowledge, with the ultimate goal of broadening children’s knowledge scope beyond story narratives that parents expect. In addition, we believe our annotation framework has the potential to be generalized to domain-specific tasks analogous to the real world that require structured external knowledge Vrandečić and Krötzsch (2014); Lehmann et al. (2015), such that clinicians use structured guidelines and knowledge for diagnosing ElSayed et al. (2023); American Diabetes Association (2011).
@@ 3 Expert Annotation Framework
In this step, experts identify an educationally suitable concept from the story text. We develop a collection of heuristics to filter candidate concepts that are tier 1 or tier 2  [^3] Tier 1 words are common and basic words. Tier 2 contains high-frequency words of various domains Beck et al. (2013). vocabulary and a concrete noun, verb, or adjective. First, we leverage the spaCy Honnibal and Montani (2017) English model to filter auxiliary words and punctuation 444tagged by ‘auxiliary’, ‘adposition’, ‘determiner’, ‘particle’, ‘punctuation’, ‘symbol’, and ‘other’ from the original story text. Then, we use AllenNLP’s Gardner et al. (2017) semantic role labeling tool to tag the latent structure of each sentence in the story context. This process identifies and retains key elements represented by semantic roles, including agents, goals, and results, which are subsequently treated as potential candidate concepts. We design the UI for step 1 to display one story section and allow experts to select highlighted candidate concepts in grey, as shown in Figure 5.
@@
This step allows experts to select real-world knowledge based on the concept selected previously. Inspired by  Xu et al. (2020)’s work of combining and filtering knowledge from Wiktionary  [^5] https://www.wiktionary.org/ and ConceptNet Speer et al. (2017) for commonsense question answering, we implement a knowledge matching module that can retrieve and rank external knowledge associated with each concept selected by the experts.
@@
We rank all retrieved triples with $1-\overline{s}+w$, where $\overline{s}$ denotes the similarity score and $w$ denotes the weight of a triple provided by ConceptNet, reflecting the combined influence and credibility of the triple by summing up the weights coming from all the sources that support it. The top six ranked triples are shown to annotators to balance between providing a sufficient selection and avoiding excessive distractions during the annotation task. We also retrieve the explanation for expert-selected concepts from Wiktionary to better facilitate experts’ annotations. The UI for this step is designed to display a list of recommended real-world knowledge triples along with the Wiktionary explanation to facilitate experts’ selection (Figure 6).
@@ 4 StorySparkQA
StorySparkQA aims to facilitate parents’ interactive story reading with appropriate real-world knowledge: practical, factual, everyday information that helps preschoolers understand the world around them. Our dataset consists of $5,868$ QA pairs annotated by children education experts leveraging our designed annotation framework. We present the core statistics of StorySparkQA in Table 4 and show one example in Figure 1.
@@ 4.2 Annotation Process
Following our annotation framework, we recruit 11 children education experts for the annotation task. The experts all have a minimum of 3 years of practical experience (e.g., kindergarten teachers) in learning science and possess relevant educational backgrounds. For each story section, experts are asked to first identify a concept from the story. Educational experts should select concepts that are most beneficial for children’s education from story text. The experts then proceed to select a real-world knowledge triple associated with the selected concept and create a QA pair based on the selected triple. In this process, experts are asked to take into account children’s cognitive and knowledge levels and write QA pairs that are most appropriate for 3- to 6-year-olds. We collect experts’ mental procedures by recording their selected concepts, real-world knowledge triples, and created QA pairs during the annotation process.
@@ 4.3 Statistics and Analysis of StorySparkQA
According to experts’ annotations, real-world knowledge relation “is a” and questions start with “what” have a much higher proportion than the others. Moreover, of all experts’ annotations, the selected concept words are mostly nouns (65.06%) and adjectives (22.26%), which are easy for children to comprehend. Considering the cognitive development of children, especially those aged 3 to 6, they are typically in the stage of language development and exploration, full of curiosity about the world Chouinard et al. (2007); Jirout and Klahr (2012). It is therefore natural for them to ask questions as a way to satisfy this curiosity. Consequently, parents are more inclined to use “what” questions with simple vocabulary to inspire children’s thinking and encourage them to actively acquire knowledge Taylor et al. (1994); Yu et al. (2019). Consistent with the actual habits of parents, experts’ annotated questions have a high consensus that “what” questions are more in line with children’s learning and cognitive characteristics.
@@
Compared with existing QA datasets for children education Xu et al. (2022); Zhao et al. (2023), StorySparkQA is unique mainly in the annotation process and data composition. Going beyond direct QA pair annotation utilized by most QA datasets (e.g., FairytaleQA and StoryQA), StorySparkQA’s annotation process provides step-by-step support with structured real-world knowledge to ease the effort that the experts need to craft children-appropriate and knowledgeable QA pairs. In addition, StorySparkQA includes not only QA pairs, but also corresponding story texts and expert-selected real-world knowledge triples, which reflects experts’ mental procedures when creating these children-appropriated, knowledgeable QA pairs and leads to a more comprehensive dataset for children’s knowledge expansion. Catering to parents’ practical needs for knowledge expansion in interactive story reading, StorySparkQA complements existing QA datasets, which often focus on narrative comprehension and commonsense question-answering. The key properties of StorySparkQA and related children education datasets are illustrated in Table 6 in Appendix A.2.
@@ 5 Benchmark Experiment
We benchmark the quality and usability of our StorySparkQA on the QAG task, which is required to meet the needs of parents to guide children to learn some real-work knowledge during practical interactive story reading, as well as existing work of developing AI-assisted storytelling and reading systems Yao et al. (2021); Dietz et al. (2021); Zhang et al. (2022). We conduct an automated evaluation, reported in Section 5.1 to measure the semantic similarity of generated QA pairs with experts-annotated QA pairs, benchmarked with a T5-Large model fine-tuned on StorySparkQA and a set of robust LLMs. Considering the limitation of automated evaluation in evaluating the educational appropriateness of generated QA pairs, we further conduct a human evaluation, reported in Section 5.2, with children education experts.
@@ 5.2 Human Evaluation
Note: * denotes p-value <0.05, and ** denotes p-value <0.01. Both of the denotations indicate results are statistically significant.
Table 3: The paired sample t-test result of children education experts in comparison of GPT-4 and T5-Large fine-tuned on StorySparkQA in the QAG task. Bolded numbers are the best scores within each dimension excluding human experts’ annotations.
@@ 5.2.1 Results and Analysis
Table 5.2 illustrates the average scores of each dimension and paired sample t-test results. We observe that expert-created QA pairs outperform those generated by models in all four dimensions. The paired sample t-test results show that experts’ annotations are significantly different in three out of four dimensions compared with models’ generation. These justify StorySparkQA’s utility in catering to parents’ real-world needs in interactive story reading.
@@ 5.3 Discussion
Comparing the best-performing SoTA LLMs in the QAG pipeline with the corresponding fine-tuned T5-Large, we can observe that the T5-Large can reliably generate QA pairs aligned more with experts’ annotations in terms of Rouge-L score according to system evaluation, regardless of whether generating QA pairs along real-world knowledge triples. Drawing from the results of our human evaluation, the fine-tuned T5-Large exhibits better capabilities in generating QA pairs that suit parents’ real-world educational expectations of interactive story reading: originating from the story and embodying educational-appropriate real-world knowledge. Worth mentioning that T5-Large only consists of $770$ million parameters, whereas Alpaca-7B, Mistral-7B, and Llama 2 in our experiments consist of $7$ billion parameters ($10$ times larger).
This observation justifies StorySparkQA’s utility in training a task-specific model that caters to parents’ real-world story reading needs on the one hand, and demonstrates the usefulness of combining structured real-world knowledge and free-form narratives in domain-specific tasks such as interactive story reading.
@@ 6 Conclusion and Future Work
In summary, we propose StorySparkQA, an expert-annotated, external-knowledge-enriched QA dataset for children education, by leveraging a novel annotation framework to facilitate scalable expert annotations through structured external knowledge. We demonstrate the effectiveness of StorySparkQA through an automated evaluation on various LLMs of generating QA pairs catering to parents’ needs and a human evaluation with children education experts.
One possible future work is refining the QAG pipeline structures and exploiting LLMs to generate QA pairs that align more closely with parents’ practical needs. Another future direction involves using StorySparkQA and language models to develop a human-AI collaborative education system (e.g., an interactive story reading system), aiding parents and educators to formulate personalized questions during story readings, while addressing their language, knowledge, or time constraints. Also, fine-tuning LLMs (e.g., Llama 3) may lead to better performance on the QAG task, which offers a future direction to refine models’ capabilities in real-world tasks like children’s education.
7 Acknowledgement
We thank the Lab of Artificial Intelligence for Education, East China Normal University, Shanghai Institute of Artificial Intelligence for Education, East China Normal University, and Shanghai Science and Technology Innovation Action Plan project (No. 21511104500) for funding this research.
@@ A.8 LLM Prompts
To utilize LLMs’ strong reasoning and generation capability as well as control GPT-generated questions as much as possible to meet the needs of parents, we carefully design our prompts.

### v3

@@ Abstract
Interactive story reading is common in early childhood education, where teachers expect to teach both language skills and real-world knowledge beyond the story. While many story reading systems have been developed for this activity, they often fail to infuse real-world knowledge into the conversation. This limitation can be attributed to the existing question-answering (QA) datasets used for children’s education, upon which the systems are built, failing to capture the nuances of how education experts think when conducting interactive story reading activities. To bridge this gap, we design an annotation framework, empowered by existing knowledge graph to capture experts’ annotations and thinking process, and leverage this framework to construct StorySparkQA dataset, which comprises $5,868$ expert-annotated QA pairs with real-world knowledge. We conduct automated and human expert evaluations across various QA pair generation settings to demonstrate that our StorySparkQA can effectively support models in generating QA pairs that target real-world knowledge beyond story content. StorySparkQA  [^1] The StorySparkQA dataset and annotation framework can also be found in: https://github.com/neuhai/StorySparkQA is available at https://huggingface.co/datasets/NEU-HAI/StorySparkQA.
@@
Jiaju Chen* East China Normal University                        Yuxuan Lu Northeastern University                        Shao Zhangthanks: Work done when visiting Northeastern University. Shanghai Jiao Tong University
@@
Yuling Sunthanks:  Corresponding Author: ylsun@cs.ecnu.edu.cn. Lab of Artificial Intelligence for Education East China Normal University
@@ 1 Introduction
Interactive story reading is common in early childhood education, where teachers often sit together with preschool children, read storybooks, and proactively engage in question-answering (QA) conversations with them Wright (1995); Isbell et al. (2004). Such guided conversations are typically grounded in but beyond the story narratives Kotaman (2013), with teachers’ expectations of guiding children to learn real-world knowledge and improving their historical, cultural, and emotional awareness Sun et al. (2024). This immersive story-based interaction has been proven to be effective in better supporting preschooler knowledge learning Zhang et al. (2024), enhancing their reading comprehension capabilities Xu et al. (2021), etc.
Despite the benefits of interactive story reading, teachers often struggle to appropriately conduct such interactive story reading with children because of multi-facet difficulties Golinkoff et al. (2019); Sun et al. (2024). Specifically, such interactive story reading needs teachers to identify the knowledge of interest during storytelling, formulate the real-world knowledge piece they want to teach in mind ("what to ask"), then ask an engaging question ("how to ask") to children at the appropriate time ("when to ask"). In home settings, most parents also lack the educational expertise necessary to guide such educational conversations Golinkoff et al. (2019); Sun et al. (2024). Meanwhile, today’s parents often hardly maintain constant focus on their children due to the need to deal with other work and family chores at the same time Zhang et al. (2022); Sun et al. (2024).
Recently, AI-assisted storytelling systems (e.g. StoryBuddy Zhang et al. (2022), TaleMate Vargas-Diaz et al. (2023), MatheMyths Zhang et al. (2024)), have demonstrated utility in children’s storytelling scenarios Dietz et al. (2021). These systems primarily utilize the verbal communication interface and advanced language models to support natural conversation with humans Mahmood et al. (2023); Chan et al. (2024); Yang et al. (2024); Dietz et al. (2021). Nevertheless, existing AI-assisted storytelling systems are not without limitations. Particularly, building on top of data resources with mostly extractive QA pairs (e.g., FairytaleQA Xu et al. (2022)) – where the answers can be found directly in the story narrative – these systems fall short in helping teaching real-world knowledge beyond the story narrative Yao et al. (2021); Zhao et al. (2022), which is one of the main expectations of parents and teachers Sun et al. (2024).
@@
StorySparkQA can benefit different research directions in the children education, particularly in better understanding domain experts’ thinking process, and training models to generate story-based QA pairs infused with real-world knowledge, with the ultimate goal of broadening children’s knowledge scope beyond story narratives that parents and teachers expect. In addition, we believe our annotation framework has the potential to be generalized to domain-specific tasks analogous to the real world that require structured external knowledge Vrandečić and Krötzsch (2014); Lehmann et al. (2015), such that clinicians use structured guidelines and knowledge for diagnosing ElSayed et al. (2023); American Diabetes Association (2011).
@@ 3 Expert Annotation Framework
In this step (UI shown in Figure 5), experts identify an educationally suitable concept from the story text. We develop a collection of heuristics to filter candidate concepts that are tier 1 or tier 2  [^3] Tier 1 words are common and basic words. Tier 2 contains high-frequency words of various domains Beck et al. (2013). vocabulary and a concrete noun, verb, or adjective. First, we leverage the spaCy Honnibal and Montani (2017) English model to filter auxiliary words and punctuation 444tagged by ‘auxiliary’, ‘adposition’, ‘determiner’, ‘particle’, ‘punctuation’, ‘symbol’, and ‘other’ from the original story text. Then, we use AllenNLP’s Gardner et al. (2017) semantic role labeling tool to tag the latent structure of each sentence in the story context. This process identifies and retains key elements represented by semantic roles, including agents, goals, and results, which are subsequently treated as potential candidate concepts.
@@
This step (UI shown in Figure 6) allows experts to select real-world knowledge based on the concept selected previously. Inspired by  Xu et al. (2020)’s work of combining and filtering knowledge from Wiktionary  [^5] https://www.wiktionary.org/ and ConceptNet Speer et al. (2017) for commonsense question answering, we implement a knowledge matching module that can retrieve and rank external knowledge associated with each concept selected by the experts.
@@
We rank all retrieved triples with $1-\overline{s}+w$, where $\overline{s}$ denotes the similarity score and $w$ denotes the weight of a triple provided by ConceptNet, reflecting the combined influence and credibility of the triple by summing up the weights coming from all the sources that support it. The top six ranked triples are shown to annotators to balance between providing a sufficient selection and avoiding excessive distractions during annotation. We also retrieve the explanation for concepts from Wiktionary to better facilitate experts’ annotations.
@@ 4 StorySparkQA
StorySparkQA aims to facilitate teachers’ interactive story reading with appropriate real-world knowledge: practical, factual, everyday information that helps preschoolers understand the world around them. Our dataset consists of $5,868$ QA pairs annotated by children education experts leveraging our designed annotation framework. We present the core statistics of StorySparkQA in Table 4 and show one example in Figure 1.
@@ 4.2 Annotation Process
Following our annotation framework, we recruit 11 children education experts for the annotation task. The experts all have a minimum of 3 years of practical experience (e.g., kindergarten teachers) in learning science and possess relevant educational backgrounds. For each story section, experts are asked to first identify a concept from the story by selecting concepts that are most beneficial for children’s education from story text. The experts then proceed to select a real-world knowledge triple associated with the selected concept and create a QA pair based on the selected triple. In this process, experts are asked to consider children’s cognitive and knowledge levels and write QA pairs that are most appropriate for 3- to 6-year-olds. We collect experts’ mental procedures by recording their selected concepts, real-world knowledge triples, and created QA pairs during the annotation process.
@@ 4.3 Statistics and Analysis of StorySparkQA
According to experts’ annotations, real-world knowledge relation “is a” and questions start with “what” have a much higher proportion than the others. Moreover, of all experts’ annotations, the selected concept words are mostly nouns (65.06%) and adjectives (22.26%), which are easy for children to comprehend. Considering the cognitive development of children, especially those aged 3 to 6, they are typically in the stage of language development and exploration, full of curiosity about the world Chouinard et al. (2007); Jirout and Klahr (2012). It is therefore natural for them to ask questions as a way to satisfy this curiosity. Consequently, teachers are more inclined to use “what” questions with simple vocabulary to inspire children’s thinking and encourage them to actively acquire knowledge Taylor et al. (1994); Yu et al. (2019). Consistent with the actual habits of teachers, experts’ annotated questions have a high consensus that “what” questions are more aligned with children’s cognitive characteristics.
@@
Compared with existing QA datasets for children education Xu et al. (2022); Zhao et al. (2023), StorySparkQA is unique mainly in the annotation process and data composition. Going beyond direct QA pair annotation utilized by most QA datasets (e.g., FairytaleQA and StoryQA), StorySparkQA’s annotation process provides step-by-step support with structured real-world knowledge to ease the effort that the experts need to craft children-appropriate and knowledgeable QA pairs. In addition, StorySparkQA includes not only QA pairs, but also corresponding story texts and expert-selected real-world knowledge triples, which reflects experts’ mental procedures when creating these children-appropriated, knowledgeable QA pairs and leads to a more comprehensive dataset for children’s knowledge expansion. Catering to teachers’ practical needs for knowledge expansion in interactive story reading, StorySparkQA complements existing QA datasets, which often focus on narrative comprehension and commonsense question-answering. The key properties of StorySparkQA and related children education datasets are illustrated in Table 6 in Appendix A.2.
@@ 5 Benchmark Experiment
We benchmark the quality and usability of our StorySparkQA on the QAG task, which is required to meet the needs of teachers to guide children to learn some real-work knowledge during practical interactive story reading, as well as existing work of developing AI-assisted storytelling and reading systems Yao et al. (2021); Dietz et al. (2021); Zhang et al. (2022). We conduct an automated evaluation, reported in Section 5.1 to measure the semantic similarity of generated QA pairs with experts-annotated QA pairs, benchmarked with a T5-Large model fine-tuned on StorySparkQA and a set of robust LLMs. Considering the limitation of automated evaluation in evaluating the educational appropriateness of generated QA pairs, we further conduct a human evaluation, reported in Section 5.2, with children education experts.
@@ 5.2 Human Evaluation
Table 3: The paired sample t-test result of children education experts in comparison of GPT-4 and T5-Large fine-tuned on StorySparkQA in the QAG task. Bolded numbers are the best scores within each dimension excluding human experts’ annotations. * means p-value <0.05, and ** means p-value <0.01, both are statistically significant.
@@ 5.2.1 Results and Analysis
Table 5.2 illustrates the average scores of each dimension and paired sample t-test results. We observe that expert-created QA pairs outperform those generated by models in all four dimensions. The paired sample t-test results show that experts’ annotations are significantly different in three out of four dimensions compared with models’ generation. These justify StorySparkQA’s utility in catering to teachers’ real-world needs in interactive story reading.
@@ 5.3 Discussion
Comparing the best-performing SoTA LLMs in the QAG pipeline with the corresponding fine-tuned T5-Large, we can observe that the T5-Large can reliably generate QA pairs aligned more with experts’ annotations in terms of Rouge-L score according to system evaluation, regardless of whether generating QA pairs along real-world knowledge triples. Drawing from the results of our human evaluation, the fine-tuned T5-Large exhibits better capabilities in generating QA pairs that suit teachers’ real-world educational expectations of interactive story reading: originating from the story and embodying educational-appropriate real-world knowledge. Worth mentioning that T5-Large only consists of $770$ million parameters, whereas Alpaca-7B, Mistral-7B, and Llama 2 in our experiments consist of $7$ billion parameters ($10$ times larger).
This observation justifies StorySparkQA’s utility in training a task-specific model that caters to teachers’ real-world story reading needs on the one hand, and demonstrates the usefulness of combining structured real-world knowledge and free-form narratives in domain-specific tasks such as interactive story reading.
@@ 6 Conclusion and Future Work
In summary, we propose StorySparkQA, an expert-annotated, external-knowledge-enriched QA dataset for children education, by leveraging a novel annotation framework to facilitate scalable expert annotations through structured external knowledge. We demonstrate the effectiveness of StorySparkQA through an automated evaluation on various LLMs of generating QA pairs catering to teachers’ needs and a human evaluation with children education experts.
One possible future work is refining the QAG pipeline structures and exploiting LLMs to generate QA pairs that align more closely with teachers’ practical needs. Another future direction involves using StorySparkQA and language models to develop a human-AI collaborative education system (e.g., an interactive story reading system) Wang et al. (2020, 2019), aiding parents and educators to formulate personalized questions during story readings, while addressing their language, knowledge, or time constraints. Also, fine-tuning LLMs (e.g., Llama 3) may lead to better performance on the QAG task, which offers a future direction to refine models’ capabilities in real-world tasks like children’s education.
@@ A.8 LLM Prompts
To utilize LLMs’ strong reasoning and generation capability as well as control GPT-generated questions as much as possible to meet the needs of teachers, we carefully design our prompts.